(8) G) A 50.0 mL sample of a gas was found to have a pressure of l.0 atm at 25 C. What will be the new volume if the temperature is changed to 50 C at constant pressure (ii) A 2.00 Lsample of Nitrogen gas at 25.0 c was found to have a pressure of 100 atm. How many moles of gas are present and what is the mass of the gas in grams? (9) The total pressure in a mixture ofNitrogen and chlorine gas was found to be 1.65 atm. if the partial pressure of Nitrogen was 0.65 atm, what is the partial pressure of Chlorine gas. (10) octane, CsH s, is a flammable hydrocarbon. It undergoes a combustion reaction (with oxygen gas) to produce Carbon dioxide and Water. (a) Write a balanced chemical equation for this reaction (b) How many mole of oxygen is consumed if 0.050 mol of octane undergoes Combustion?
Answer #1

Q8.

V = 50 mL, P = 1 atm, T = 25°C = 298K

find new volume when T = 50°C = 323 K

so

V1/T1 = V2/T2

substitute

50/298 = V2/323

V2 = 323/298*50 = 51.20 mL

ii

V = 2 L of N2gas T = 298K, P = 1 atm

find moles of gas present

mol from ideal gas law

PV = nRT

n = PV/(RT)

n = (1)(2)/(0.082*298) = 0.08184645 moles of gas

mass = mol*MW

MNW of N2 = 28 g/mol

mass = mol*MW = (0.08184645*28) = 2.2917006 g of N2